I have to solve by the addition method. I know that I am supposed to mutiply both sides by an integer and then solve. I am coming up with the wrong answer
3x+5y=-11
x-2y=11
I keep getting 15 for x

Answer by stanbon(75887)

3x+5y=-11
x-2y=11
------------
Multiply thru the 2nd equation by -3:
3x+5y = -11
-3x+6y = -33
--------------------
Add the equations:
11y = -44
---
y = -4
------
Sustitute for "y" in x-2y = 11 and solve for "x".
x -2(-4) = 11
x+8 = 11
x = 3
=====================
Cheers,
Stan H.